OpenStudy (anonymous):

The table below shows the number of wafers in different numbers of packs: Number of Packs Number of Wafers 2 16 3 24 4 32 5 40 Is this true or false? The numbers in the table represent a proportional relationship. True False

Mehek (mehek14):

what do you think?

OpenStudy (anonymous):

True

Mehek (mehek14):

check \(\dfrac{2}{16}=\dfrac{1}{8}\\\dfrac{3}{24}=\dfrac{1}{8}\)

Mehek (mehek14):

you are correct

OpenStudy (anonymous):

Yay I thought it was true but I wan't sure!

Mehek (mehek14):

just write them down as fractions and if they equal then it is proportional